# Benchmarking ECG Foundational Models: A Reality Check Across Clinical Tasks

### v1

@@ Abstract
The 12-lead electrocardiogram (ECG) is a long-standing diagnostic tool. Yet machine learning for ECG interpretation remains fragmented, often limited to narrow tasks or datasets. Foundation models promise broader adaptability, but their generalization across diverse ECG tasks is not well understood. We benchmarked eight ECG foundation models on 26 clinically relevant tasks using 12 public datasets comprising 1,650 regression and classification targets. Models were evaluated under fine-tuning and frozen settings, with scaling analyses across dataset sizes. Results show heterogeneous performance across domains: in the most widely studied domain, adult ECG interpretation, three foundation models consistently outperformed strong supervised baselines. In contrast, ECG-CPC, a compact structured state-space model pretrained on HEEDB, dominated other categories where most foundation models failed to surpass supervised learning. Foundation models also displayed distinct scaling behaviors with dataset size, which are critical for small-scale clinical applications. Overall, while foundation models show promise for adult ECG analysis, substantial gaps remain in cardiac structure, outcome prediction, and patient characterization. Notably, ECG-CPC’s strong performance despite being orders of magnitude smaller and consuming minimal computational resources highlights untapped opportunities for advancing ECG foundation models.
Figure 1: Overview of the benchmarking pipeline for ECG foundation models.
@@ 1 Introduction
Clinical relevance Electrocardiography (ECG) is a widely used non-invasive tool for assessing cardiac function and systemic physiology (Siontis et al., 2021). Accurate interpretation is essential for detecting myocardial infarctions (Strodthoff et al., 2020), evaluating cardiovascular risk (Bhatia & Dorian, 2018), and guiding clinical decisions (Rokos et al., 2010). ECG features also reflect systemic conditions such as electrolyte imbalances (Diercks et al., 2004), metabolic disorders (Wald, 2006), and physiological factors (Lopez Alcaraz et al., 2025), broadening its clinical utility. The growing availability of large-scale ECG datasets (Wagner et al., 2022; Gow et al., 2023; Ribeiro et al., 2021) has made machine learning increasingly pivotal for automated interpretation, enabling population-scale screening (Kalmady et al., 2024) and supporting clinical workflows (Graf et al., 2024).
@@
Promise of foundational models for ECG Foundational models (FMs) have transformed fields such as natural language processing (Myers et al., 2024) and computer vision (Awais et al., 2025), where large-scale pretraining produces robust and transferable representations across tasks and domains (Subramanian et al., 2023). This approach is now emerging for biomedical time series, with ECG as a natural testbed, given its ubiquity and clinical relevance. Clinical FMs offer three key advantages: (i) higher predictive performance than training from scratch, (ii) improved label efficiency via pretrained representations, and (iii) the utility of FMs as frozen feature extractors for downstream tasks. These benefits could enable stronger predictive models, support rare-disease studies on limited data, and guide the choice of foundation models for specific applications.
Research gap Despite their promise, foundational models for ECG are under-evaluated. Prior studies often only compare FMs against weak baselines (Li et al., 2025; Kim, 2024; Na et al., 2024; Liu et al., 2024), and benchmarks are limited in scope, focusing on narrow datasets or tasks and hindering generalizable conclusions (Na et al., 2024; Liu et al., 2024). Key aspects such as label efficiency, cross-domain representation quality, and trade-offs between fine-tuning and frozen evaluation remain poorly understood (Li et al., 2025; Kim, 2024; Na et al., 2024; Liu et al., 2024). A systematic, like-for-like comparison of available FMs, leveraging a unified codebase, is still lacking.
Contributions of this work We present a comprehensive benchmark of ECG FMs, covering classification and regression across heterogeneous clinical open-source datasets. Our contributions are threefold: (1) FM comparison: We evaluate FMs in different evaluation modes against strong supervised baselines such as structured state-space (S4) models. While FMs often surpass weak baselines, only the strongest match S4 performance. FMs reveal heterogeneous, highly task-specific results. Certain FMs can effectively serve as frozen feature extractors for certain tasks with competitive performance. (2) Scaling: We investigate the prospects of finetuning FMs as a possible solution to the challenges of small sample sizes. Strong FMs can indeed increase label efficiency by factors of 2.5-9 compared to a strong supervised baseline. (3) FM training: We propose ECG-CPC, a lightweight foundation model using structured state-space models as the architecture and trained with minimal computational resources, which turns out as one of the strongest models in the benchmark.
@@ 2 Background
AI-enhanced ECG Electrocardiography (ECG) is a widely used, non-invasive tool for assessing cardiac function and systemic health (Siontis et al., 2021). Machine learning and deep learning now underpin automated ECG interpretation, improving arrhythmia detection, risk stratification, and clinical decision support. Models of different architectural flavors, from CNNs to RNNs to transformers, trained through supervised learning, have been shown to achieve strong performance. However, they require large, curated datasets and often generalize poorly across populations, devices, and settings (Ribeiro et al., 2020; Hannun et al., 2019; Hong et al., 2020). Benchmarking robust supervised baselines across multiple datasets remains crucial for assessing the benefits of representation learning and pretraining (Strodthoff et al., 2020; Nonaka & Seita, 2021; Hong et al., 2020).
ECG foundation models Inspired by the success of foundation models in language and vision, several approaches have been proposed for ECG. Architectures include CNNs (mainly ResNet variants), transformers (often with convolutional encoders), and structured state-space models. Pretraining methods vary from supervised and weakly supervised to contrastive and non-contrastive self-supervision. Pretraining datasets also differ widely, from Computing in Cardiology 2021 challenge subsets (Reyna et al., 2021b) to MIMIC-IV ECG (Gow et al., 2023) to the large-scale HEEDB dataset (Koscova et al., 2024).
Table 1: Summary of the eight ECG foundation models and two supervised baselines evaluated in this work, including their backbone architectures, training paradigms, pretraining datasets and sizes, and parameter sizes. Pretraining datasets: a: HEEDB; b: Chapman, Ningbo, CODE-15%; c: MIMIC-IV-ECG; d: CODE, PTB, CPSC2018, CPSC-Extra, PTB-XL, Georgia, Chapman, Ningbo, Hefei, SPH, MIMIC-IV-ECG; e: CPSC2018, CPSC-Extra, PTB-XL, Georgia, Ningbo, Chapman, MIMIC-IV-ECG
@@
ST-MEM Transformer MAE 174kb 90.3M
MERL Transformer Weak sup., contrastive 800kc 4.6M
@@
ECG-FM Transformer MLM+contrastive 2.5Me 93.9M
ECG-CPC SSM CPC 10.7Ma 3.8M
Net-1D CNN - - 33.8M
S4 SSM - - 2.2M
Benchmarking foundation models The benchmarking of foundation models has a long tradition in other fields such as computer vision (Goldblum et al., 2023) and NLP (Chang et al., 2024). In the medical domain, most efforts have so far focused on medical imaging (Neidlinger et al., 2024; Lee et al., 2025). Recently, benchmarking results for EEG foundation models have been put forward (Xiong et al., 2025). However, no large-scale, comprehensive benchmark for ECG foundation models exists.
@@ 3.1 Models
Table 1 summarizes the investigated models, including backbones, pretraining methods, and datasets, and parameter counts. As custom pretraining is infeasible, only foundation models with publicly available pretrained weights are included and integrated via wrapper modules into a common evaluation framework. Unfortunately, the lack of public weights excludes many proposed ECG FMs. We benchmark eight foundational models and two supervised baseline models as described below.
Foundation models We consider diverse architectures and pretraining strategies. ECGFounder (Li et al., 2025) uses a RegNet-inspired CNN pretrained on HEEDB with a supervised loss. ECG-JEPA (Kim, 2024) employs a transformer-based joint embedding predictive architecture (JEPA) (Assran et al., 2023). We use the multi-block variant. ST-MEM (Na et al., 2024) uses a ViT-1D transformer trained as a masked autoencoder (MAE) (He et al., 2022). MERL (Liu et al., 2024) applies contrastive text-signal alignment (Radford et al., 2021), with the ResNet18 variant evaluated. ECGFM-KED (Tian et al., 2024) uses a ResNet backbone and minimizes a contrastive loss between signals and ECG report text. HuBERT-ECG and ECG-FM (McKeen et al., 2024) are transformer-based, pretrained via masked language modeling, with ECG-FM additionally incorporating a sequence-level contrastive loss. Finally, we pretrain a structured state-space (SSM) model on HEEDB using contrastive predictive coding (CPC) (van den Oord et al., 2018; Mehari & Strodthoff, 2023) (see Supplementary Material).
Supervised baselines We evaluate two backbones trained from scratch: Net-1D, from ECGFounder (Li et al., 2025), and S4 (Mehari & Strodthoff, 2023), a structured state-space model known for its ability to capture long-range dependencies (Gu et al., 2022). S4 has proven to be a strong baseline on PTB-XL and other datasets (Mehari & Strodthoff, 2023; Strodthoff et al., 2024).
@@ 3.2 Datasets and benchmarking tasks
We group the benchmark datasets into seven categories based on the qualitative nature of the underlying labels. See Table 3 in the Supplementary Material for details on sample sizes and task characteristics. Adult/pediatric ECG interpretation: These two categories, the most studied in the literature, involve predicting diagnostic ECG statements from cardiologists. Benchmarks include PTB (Bousseljot et al., 1995), Ningbo (Reyna et al., 2022), CPSC2018 and CPSC-Extra (Reyna et al., 2021a; 2022), Georgia (Reyna et al., 2021a; 2022), Chapman (Reyna et al., 2021a; 2022), SPH (Liu et al., 2022a; b), CODE-15% (Ribeiro et al., 2021), and PTB-XL (Wagner et al., 2022; 2020). A separate category includes ZZU pECG (Tan et al., 2025; Jian et al., 2025) for pediatric ECG interpretation. Cardiac structure and function: This category predicts outcomes from complementary modalities, here echocardiography. We use the EchoNext dataset (Elias & Finer, 2025), which is used to capture cardiac structure and function. Cardiac and non-cardiac outcomes: These categories cover the prediction of cardiac and non-cardiac discharge diagnoses from the first emergency department ECG (Strodthoff et al., 2024), distinguishing cardiac (ICD-10 chapter I) and non-cardiac diagnoses. Acute care predictions: ECG is a key modality for acute care decisions. We predict clinical deterioration, mortality at multiple time frames, and ICU admission (Alcaraz et al., 2025), training a joint model for cardiac, non-cardiac, and acute care outcomes to save computational resources. Patient characteristics: This category includes tasks where ECGs predict non-diagnostic patient characteristics, such as sex, age, biometrics, ECG features, laboratory values (Alcaraz & Strodthoff, 2024), and vital signs (Gow et al., 2023). The datasets span small specialized cohorts to large population studies, covering diverse classification and regression tasks. We train a single model for all tasks in this category using combined classification and regression losses.
@@ 3.3 Methodology
Finetuning/linear evaluation Pretrained models are supplemented with a linear classification head matching the downstream task outputs. We optimize binary cross-entropy or mean absolute error (MAE) using AdamW with a constant learning rate of 1e-3 and weight decay of 1e-3, performing model selection on the validation set via AUROC or MAE. In order to avoid negative interference for multiple regression targets at different scales, we use z-normalized targets (based on training set statistics). During finetuning, layer-dependent learning rates are applied: model architectures are divided into two parts, corresponding learning rates are scaled down by factors of 100 and 10 relative to that of the prediction head. Batch-norm statistics are frozen for linear evaluation and frozen evaluation. For frozen evaluation, we replace the linear prediction head with a learnable query attention head (Bardes et al., 2024) operating on the sequence of output tokens before pooling. If the model allows to adjust the input size, random 2.5s input crops are used for training (Mehari & Strodthoff, 2023), and predictions are averaged across non-overlapping sliding windows during test time.
Evaluation For classification, we report macro-averaged AUROC, as a measure of overall discriminative performance. For regression, we report the average MAE across z-normalized predictions and targets. Individual label AUROC and MAE values are provided in the Supplementary Material. Statistical uncertainty is assessed via empirical bootstrapping on the test set ($n=1,000$ iterations). Pairwise comparisons are performed by bootstrapping performance differences, if the 95% confidence interval excludes zero, the difference is considered significant. Rankings are assigned based on these comparisons, with ties indicating no significant difference, capturing both relative performance, and statistical significance across task categories.

### v2

# Benchmarking ECG FMs: A Reality Check Across Clinical Tasks

@@ Abstract
The 12-lead electrocardiogram (ECG) is a long-standing diagnostic tool. Yet machine learning for ECG interpretation remains fragmented, often limited to narrow tasks or datasets. FMs promise broader adaptability, but fundamental questions remain: Which architectures generalize best? How do models scale with limited labels? What explains performance differences across model families? We benchmarked eight ECG FMs on 26 clinically relevant tasks using 12 public datasets comprising 1,650 regression and classification targets. Models were evaluated under fine-tuning and frozen settings, with scaling analyses across dataset sizes. Results show heterogeneous performance across domains: in adult ECG interpretation, three FMs consistently outperformed strong supervised baselines. In contrast, ECG-CPC, a compact structured state-space model, dominated 5 of 7 task categories, demonstrating that architecture matters more than scale. FMs improved label efficiency 3.3-9× over supervised baselines, though scaling behaviors varied across architectures. Representation analysis reveals that models with similar performance learn markedly different internal structures, suggesting multiple viable paths to effective ECG representation. Overall, while FMs show promise for adult ECG analysis, substantial gaps remain in cardiac structure, outcome prediction, and patient characterization. ECG-CPC’s strong performance despite being orders of magnitude smaller challenges the assumption that FM quality requires massive scale, highlighting architectural inductive biases as an untapped opportunity.
Figure 1: Overview of the benchmarking pipeline for ECG FMs.
@@ 1 Introduction
Clinical relevance Electrocardiography (ECG) is a widely used non-invasive tool for assessing cardiac function and systemic physiology (Siontis et al., 2021). Accurate interpretation is essential for detecting myocardial infarctions (Strodthoff et al., 2020), evaluating cardiovascular risk (Bhatia and Dorian, 2018), and guiding clinical decisions (Rokos et al., 2010). ECG features also reflect systemic conditions such as electrolyte imbalances (Diercks et al., 2004), metabolic disorders (Wald, 2006), and physiological factors (Lopez Alcaraz et al., 2025), broadening its clinical utility. The growing availability of large-scale ECG datasets (Wagner et al., 2022; Gow et al., 2023; Ribeiro et al., 2021) has made machine learning increasingly pivotal for automated interpretation, enabling population-scale screening (Kalmady et al., 2024) and supporting clinical workflows (Graf et al., 2024).
@@
Promise of FMs for ECG We follow the original definition from the Stanford Institute for Human-Centered Artificial Intelligence (Bommasani et al., 2022) and define foundational models (FMs) as “models trained on broad data (generally using self-supervision at scale) that can be adapted to a wide range of downstream tasks”. FMs have transformed fields such as natural language processing (Myers et al., 2024) and computer vision (Awais et al., 2025), where large-scale pretraining produces robust and transferable representations across tasks and domains (Subramanian et al., 2023). FMs now also emerge in biomedical domains, including pathology (Chen et al., 2024), retinal imaging (Zhou et al., 2023), and biomedical time series, where ECG serves as a natural testbed given its ubiquity and clinical relevance. Clinical FMs offer three key advantages: (i) higher predictive performance than training from scratch, (ii) improved label efficiency via pretrained representations, and (iii) the utility of FMs as frozen feature extractors for downstream tasks. These benefits could enable stronger predictive models, support rare-disease studies on limited data, and guide the choice of FMs for specific applications.
Research gap Despite their promise, ECG FMs remain inadequately evaluated. Prior studies compare FMs against weak baselines (Li et al., 2025; Kim, 2024; Na et al., 2024; Liu et al., 2024), and existing benchmarks focus on narrow datasets or single task categories, preventing generalizable conclusions about model capabilities (Na et al., 2024; Liu et al., 2024). Critically, fundamental questions remain unanswered: Do larger FMs consistently outperform smaller, task-specific models? Which architectural choices from Transformers, CNNs, or state-space models yield the most transferable representations? How do different pretraining strategies affect label efficiency across clinical domains? Without systematic comparison under controlled conditions, the field cannot distinguish genuine advances from benchmark overfitting or favorable task selection.
Contributions of this work This work addresses three research questions central to ECG FM development: Architecture: Which backbone architectures generalize best across diverse ECG tasks?; Efficiency: How do FMs scale with labeled data compared to supervised baselines?; Representation: What explains performance differences between architecturally distinct models?
We present a comprehensive benchmark covering 8 FMs, 12 datasets, and 26 clinical tasks across classification and regression. Our key findings are: (1) Architecture matters more than scale. Structured state-space models (SSMs) outperform Transformer-based models on 5 of 7 task categories despite being much smaller. We introduce ECG-CPC, a lightweight SSM trained with minimal resources, which matches or exceeds all evaluated models, challenging the assumption that FM quality scales primarily with parameter count. (2) FMs improve label efficiency 3.3-9×. Scaling analysis shows strong models substantially reduce labeled data requirements compared to supervised baselines, especially for outcome prediction, though gains vary across architectures, guiding model selection under data constraints. (3) Similar performance arises from divergent representations. Centered kernel alignment (CKA) (Kornblith et al., 2019; Nguyen et al., 2021) analysis reveals that models with comparable downstream performance learn markedly different internal feature structures, suggesting multiple viable paths to effective ECG representation and raising questions about whether task performance alone suffices for evaluating FMs.
@@ 2 Background
AI-enhanced ECG Machine learning and deep learning now underpin automated ECG interpretation, improving arrhythmia detection, risk stratification, and clinical decision support. Models of different architectural flavors, from CNNs to RNNs to Transformers, trained through supervised learning, have been shown to achieve strong performance. However, they require large, curated datasets and often generalize poorly across populations, devices, and settings (Ribeiro et al., 2020; Hannun et al., 2019; Hong et al., 2020). Benchmarking robust supervised baselines across multiple datasets remains crucial for assessing the benefits of representation learning and pretraining (Strodthoff et al., 2020; Nonaka and Seita, 2021; Hong et al., 2020).
ECG FMs Inspired by the success of FMs in language and vision, several approaches have been proposed for ECG. Architectures include CNNs (mainly ResNet variants), transformers (often with convolutional encoders), and structured state-space models. Pretraining methods vary from supervised and weakly supervised to contrastive and non-contrastive self-supervision. Pretraining datasets also differ widely, from Computing in Cardiology 2021 challenge subsets (Reyna et al., 2021b) to MIMIC-IV-ECG (Gow et al., 2023) to the large-scale HEEDB dataset (Koscova et al., 2024).
Table 1: Summary of the eight ECG FMs and two supervised baselines evaluated in this work, including their backbone architectures, training paradigms, pretraining datasets and sizes, and parameter sizes. Pretraining datasets: a: HEEDB; b: Chapman, Ningbo, CODE-15%; c: MIMIC-IV-ECG; d: CODE, PTB, CPSC2018, CPSC-Extra, PTB-XL, Georgia, Chapman, Ningbo, Hefei, SPH, MIMIC-IV-ECG; e: CPSC2018, CPSC-Extra, PTB-XL, Georgia, Ningbo, Chapman, MIMIC-IV-ECG; * indicates models that were trained in this work, all other models were taken from the literature.
@@
ST-MEM Transformer Masked Autoencoder 174kb 90.3M
MERL CNN Weak sup., contrastive 800kc 4.6M
@@
ECG-FM Transformer MLM+contrastive 1.5Me 93.9M
ECG-CPC* SSM CPC 10.7Ma 3.8M
Net1D* CNN - - 33.8M
S4* SSM - - 2.2M
Benchmarking FMs The benchmarking of FMs has a long tradition in other fields such as computer vision (Goldblum et al., 2023) and NLP (Chang et al., 2024). In the medical domain, most efforts have so far focused on medical imaging (Neidlinger et al., 2025; Lee et al., 2025). Recently, benchmarking results for EEG FMs have been put forward (Xiong et al., 2025). However, no large-scale, comprehensive benchmark for ECG FMs exists.
@@ 3.1 Models
Table 1 summarizes the investigated models, including backbones, pretraining methods, and datasets, and parameter counts. As custom pretraining is infeasible, only FMs with publicly available pretrained weights are included and integrated via wrapper modules into a common evaluation framework. We benchmark eight FMs and two supervised baseline models as described below. Minor discrepancies in parameter counts stem from adjusting the classification head to each dataset’s label count, and for multimodal models we use only the signal encoder, focusing on ECG signals rather than full released models with text or other modalities; note that these encoders retain an implicit advantage from having been pretrained alongside additional modalities.
FMs We consider diverse architectures and pretraining strategies. ECGFounder (Li et al., 2025) uses a RegNet-inspired CNN pretrained on HEEDB with a supervised loss. ECG-JEPA (Kim, 2024) employs a transformer-based joint embedding predictive architecture (JEPA) (Assran et al., 2023). We use the multi-block variant. ST-MEM (Na et al., 2024) uses a ViT-1D transformer trained as a masked autoencoder (He et al., 2022). MERL (Liu et al., 2024) applies contrastive text-signal alignment (Radford et al., 2021), with the ResNet18 variant evaluated. ECGFM-KED (Tian et al., 2024) uses a ResNet backbone and minimizes a contrastive loss between signals and ECG report text. HuBERT-ECG and ECG-FM (McKeen et al., 2025) are transformer-based, pretrained via masked language modeling, with ECG-FM additionally incorporating a sequence-level contrastive loss. Finally, we pretrain a structured state-space (SSM) model on HEEDB using contrastive predictive coding (CPC) (van den Oord et al., 2018; Mehari and Strodthoff, 2023) (see Supplementary Material).
Supervised baselines We evaluate two backbones trained from scratch: Net1D, from ECGFounder (Li et al., 2025), and S4 (Mehari and Strodthoff, 2023), a structured state-space model known for its ability to capture long-range dependencies (Gu et al., 2022). S4 has proven to be a strong baseline on PTB-XL and other datasets (Mehari and Strodthoff, 2023; Strodthoff et al., 2024).
@@ 3.2 Datasets and benchmarking tasks
We group the benchmark datasets into seven categories based on the qualitative nature of the underlying labels. See Table 2 in the Supplementary Material for details on sample sizes and task characteristics. Adult/pediatric ECG interpretation: These two categories, the most studied in the literature, involve predicting diagnostic ECG statements from cardiologists. Benchmarks include PTB (Bousseljot et al., 1995), Ningbo (Reyna et al., 2022), CPSC2018 and CPSC-Extra (Reyna et al., 2021a; 2022), Georgia (Reyna et al., 2021a; 2022), Chapman (Reyna et al., 2021a; 2022), SPH (Liu et al., 2022a; b), CODE-15% (Ribeiro et al., 2021), and PTB-XL (Wagner et al., 2022; 2020). A separate category includes ZZU pECG (Tan et al., 2025; Jian et al., 2025) for pediatric ECG interpretation. Cardiac structure and function: This category predicts outcomes from complementary modalities, here echocardiography. We use the EchoNext dataset (Elias and Finer, 2025), which is used to capture cardiac structure and function. Cardiac and non-cardiac outcomes: These categories cover the prediction of cardiac and non-cardiac discharge diagnoses from the first emergency department ECG (Strodthoff et al., 2024), distinguishing cardiac (ICD-10 chapter I) and non-cardiac diagnoses. Acute care predictions: ECG is a key modality for acute care decisions. We predict clinical deterioration, mortality at multiple time frames, and intensive care unit (ICU) admission (Alcaraz et al., 2025), training a joint model for cardiac, non-cardiac, and acute care outcomes to save computational resources. Patient characteristics: This category includes tasks where ECGs predict non-diagnostic patient characteristics, such as sex, age, biometrics, ECG features, laboratory values (Alcaraz and Strodthoff, 2025), and vital signs (Gow et al., 2023). The datasets span small specialized cohorts to large population studies, covering diverse classification and regression tasks. We train a single model for all tasks in this category using combined classification and regression losses.
@@ 3.3 Methodology
Finetuning/linear evaluation Pretrained encoders are augmented with a linear head matching each downstream task. All models use the standard 12 ECG leads without additional resampling or filtering. Training optimizes binary cross-entropy (classification) or MAE (regression) with AdamW (learning rate $1\times 10^{-3}$, weight decay $1\times 10^{-3}$), batch size 64, for 100 epochs, with model selection on the validation set via macro AUROC or MAE. To mitigate scale interference in multi-target regression, targets are z-normalized using training-set statistics. During fine-tuning, we apply layer-dependent learning rates: model backbones are split into two parts whose learning rates are reduced by factors of 100 and 10 relative to the prediction head (see Table 33 for ablation). Batch-normalization statistics are frozen for linear and frozen evaluation. For frozen evaluation, the linear head is replaced with a learnable query-attention head (Bardes et al., 2024), which uses a single trainable query vector to attend over encoder tokens and produce a pooled representation, providing a lightweight, parameter-efficient pooling mechanism for frozen embeddings. When supported by the architecture, we use 2.5-second ECG segments instead of full 10-second recordings, as longer windows substantially increase compute with minimal performance gain. At inference, predictions are averaged across four non-overlapping 2.5-second segments, improving performance over single-segment evaluation. These choices follow established discriminative ECG practice (Mehari and Strodthoff, 2023). In line with modern robustness recommendations, we do not fix a global random seed.
Evaluation For classification, we report macro-averaged AUROC as the primary measure of overall discriminative performance. For regression, we report mean absolute error (MAE) averaged across z-normalized predictions and targets. Per-label AUROC and MAE values are provided in the Supplementary Material. Statistical uncertainty is estimated via empirical bootstrapping on the test set (n=1,000n=1{,}000). Pairwise model comparisons are conducted by bootstrapping performance differences; differences whose 95% confidence interval excludes zero are deemed significant. Rankings are derived from these tests, with ties indicating no significant difference, thereby reflecting both relative performance and statistical uncertainty across task categories. We adopt macro AUROC as the primary metric because it is the standard threshold-free ranking measure of discriminative ability. For clinical deployment, threshold-dependent metrics (e.g., sensitivity and specificity at label-specific operating points) are also relevant; an exploratory consistency analysis of rankings under such metrics is provided in Appendix A.10. Moreover, recent theoretical and empirical evidence (McDermott et al., 2024) indicates that AUROC is generally more reliable than alternatives such as AUPRC, particularly under label imbalance. Zero-shot evaluation is excluded because most ECG foundation models in our benchmark lack semantic label supervision, making zero-shot predictions neither meaningful nor comparable across architectures. Instead, consistent with current best practice for medical time-series FMs and the definition of foundation models in Bommasani et al. (2022), we focus on linear probing, frozen-feature evaluation, and full fine-tuning, which more reliably assess representation quality and adaptability.